OpenStudy (anonymous):

simplify: 36a^2/7 multiplied by 5/18a

OpenStudy (lgbasallote):

\[\frac{36a^2}{7} \times \frac{5}{18a}?\]

OpenStudy (anonymous):

yeah, thats right

OpenStudy (anonymous):

factor everything at the top and bottom. then cancel common factors from top and bottom...